The line of symmetry of the parabola whose equation is y = ax^2 - 4x + 3 is x = -2. What is the value of "a"?
A.) -2
B.) -1
C.) -1/2

Found 2 solutions by ikleyn, stanbon:
Answer by ikleyn(52778)

.
-1. Option B).



Answer by stanbon(75887)

The line of symmetry of the parabola whose equation is y = ax^2 - 4x + 3 is x = -2. What is the value of "a"?
Line of symmetry occurs where x = -b/(2a)
Solve for "a":: -b/2a = 4/(2a) = -2
-4a = 4
a = -1
------------
Cheers,
Stan H.
----------------
A.) -2
B.) -1
C.) -1/2